OpenStudy (anonymous):

Write the equation of the line that passes through (3, 6) and (0, –3) in slope-intercept form. y = 3x – 6 y = 3x – 3 y = 3x + 3 y = 3x + 6

OpenStudy (radar):

One thing for sure if the choices are correct, the slope m is a positive 3. As the coefficient of x in all 4 choices is a three.

OpenStudy (radar):

In addition, one of the points is the y intercept, when x = 0, y is -3. My first guess is the second one. ..........But lets work it out.

OpenStudy (radar):

Slope m is delta y divided by delta x. Delta means the change in x or y between the two points. delta y= 6-(-3) = 9 delta x=3 - 0 = 3 9/3 = 3, just as we thought!

OpenStudy (radar):

now using just one of the points we can develop our slope-intercept formula.

OpenStudy (radar):

using the first given point (we could of used the other one= the choice is arbitrary. 6 - y 3=------ 3- x solving 9-3x=6-y y=3x+6-9 y=3x-3 Yep looks like the second one

OpenStudy (radar):